The math club consists of 9 people. How many ways can president, vice-president, and secretary be chosen in this club?
Answer by Edwin McCravy(20054)

We can choose the president any of 9 ways.

For each of those 9 ways, there are 8 remaining ways to choose the vice-president.

That's 9*8 = 72 ways to choose the president and the vice-president.

For each of those 72 ways, there are 7 remaining ways to choose the secretary.

That's 72*7=504 ways to choose the president, the vice-president, and the secretary.

Answer and how to get it:  9*8*7 = 9P3 = 504 ways.

Edwin